OpenStudy (anonymous):

How many minutes is it before 12 noon if 42 minutes ago it was five times as many minutes past 9 a.m

OpenStudy (anonymous):

lol wtf?

OpenStudy (anonymous):

what wtf this is question didnt u see it?

OpenStudy (anonymous):

spongebob

OpenStudy (anonymous):

sath help me on google doc

OpenStudy (anonymous):

mary ok

OpenStudy (anonymous):

is it x-42 = 5x? i don't understand the question.

OpenStudy (anonymous):

tryceo do youhave gmail

OpenStudy (anonymous):

sath why areigng you hellp me n gool e doc

OpenStudy (anonymous):

well....i got an answwer

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

Suppose it was x mins before 12 noon. So 9am+5x+42mins+x=12noon 6x+42mins=12pm-9am

OpenStudy (anonymous):

sath wh arent ou helpin me on google doooc

OpenStudy (anonymous):

6x=(180-42)mins x=23mins

OpenStudy (anonymous):

thus correct answer is 23 mins

OpenStudy (anonymous):

oh i see.. it was in hours... I guess that makes sense.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

stah go o google doc

OpenStudy (anonymous):

why diid oyu leva google doc